Shift summary for eve 11/11/99

Jason Hamm (jhamm@physics.arizona.edu)
Thu, 11 Nov 1999 22:56:47 -0700 (MST)


Shift Summary for eve shift, 11/11/99

On Shift: Wah, J. Hamm

Hours of beam available at all intensities: 0
Hours of beam available at nominal intensity: 0
Hours of E799 data (special runs not included): 0
Hours of special runs (laser scans, muon runs...): .75

Accounting of lost time: MI, TeV problems

========================================================================
Runs ended this shift:

14947: Laser scan

Tapes brought to FCC this shift: none
========================================================================
Accesses during shift: Valeri and Sasha G for KAMI stuff
Jason H, Wah, and Julie for bad dynode connection
========================================================================
Overview of shift: Finishing up shutdown tasks, waiting for beam
========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
16:00 AT modifies L3 code to use only hardware
clusters when computing Pt^2 for the pi0mm tag
(this change should increase the acceptance for
this mode while leaving eemumu unchanged). Also
changed namelist to do both hardware and software
clustering for triggers 5 and 6 (used to be only
software). Code remade by RSK, will be effective
from run 14943 ->

16:10 SET pluggs in 1443 cards for regenerator HV.
Discovered that card for channels 80-95 is bad.
Will replace later.

16:17 JRL reports that 4X kumquat is back in the
readout, but there's still a problem with 4X-5
upstream. Will look at during next downtime.

16:30 JRL changes rate at which T0 will update, from
every 500 events to every 100 events.

16:50 2 new bottles of argon-ethane on east side.

17:30 VJ,SG,KN finish access to install aerogel detector
prototype for KAMI beam test. Also checked -6V
supply on crate 52 (had gotten error from SDAQ) -
looks ok. Also adjusted VV' HV for a few channels
to improve efficiency.

Start cosmic run, notice missing HCC bit in laser
triggers (1382). JCH and Wah make access to
examine dynode signal into ET; lack of one points
to blockhouse problem. Page Julie. She tweaked
the dynode connection in blockhouse, which
restored the bit.

21:00 Channel 13: TeV B2 power supply problem, MI
ground fault. Est. 2-4 hours for beam.

22:00 PSh requests another laser scan. Start run 14947.

22:45 End run 14947 - good scan.

========================================================================
Follow up Notes on Previous Shift
========================================================================
Items to Follow up on Next Shift

When beam returns:

- Notify Bob T.
- Take muon run
- Flip NM4AN polarity to +
- Take pedestal run
- Update constants

========================================================================